OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −3(2u − 1) −5u − 3 −6u + 3 −5u + 3 −6u − 3

OpenStudy (anonymous):

D Right?

OpenStudy (anonymous):

no

OpenStudy (jackellyn):

Show your work, if you have any mistakes we can correct you.

OpenStudy (anonymous):

-3 x2u = -3 x -1=

OpenStudy (king):

@Deldrickg when u multiply -3,-1 -3*-1=(3*1)(-1*-1) =3*1 =3 so -6u+3

OpenStudy (king):

B

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

negative and a negative = positive :)

OpenStudy (anonymous):

but only when multiplying